three more than five times a number is the same as thirty more than twice the number. find the number
Answer by ewatrrr(24785)


Hi
three more than five times a number is the same as thirty more than twice the number.
Let x represent the number
Question states***
5x + 3 = 2x + 30
Solving for x
3c = 27
c = 9
CHECKING our Answer***
45+3 = 18+30 = 48